FEL Upgrade Project Weekly Brief
January 13-17, 2003

Commissioning, hardware qualification, and installation activities in the FEL Upgrade continued this week with the
first successful preparation of a photoactive cathode, first high voltage on the gun, installation of vacuum qualified
chambers in the optical chicane, and installation of sextupole magnets in both arcs.

Note Change in Date (Again) for 2003 FEL Users/LPC Meeting and FEL Science Workshop:
March 19-21, 2003.

Because of scheduling conflicts with key members of the external steering committee for the March FEL Science
Workshop, the dates for the next FEL Users/LPC Meeting (now March 19) and the subsequent FEL Science
Workshop (now March 20-21) have been changed.

Commissioning Progress:
Details are given in the WBS sections below. A summary is as follows:
1)  We successfully processed the GaAs cathode and obtained sufficient quantum yields to proceed with high
       voltage processing. The gun HVPS and SF6 enclosures were successfully mated up to the gun for operation and high
      voltage processing began on Thursday night.  Before the processing was to restart this morning (Friday), we
      observed a crack in the GaAs cathode that will necessitate changing out the cathode and rebaking the gun
      section (but not the entire injector assembly).  Despite this minor setback we are pleased that the gun system
      passed a fully integrated test.
2)  Two of the four vacuum qualified chambers were installed in the optical chicane bypass magnets (DW's)
3)  The vacuum hardware in the reinjection portion of the machine arc (Region 5F) was installed this week except
for one small (2 inch ) coupling.
4) The set of initial sextupole magnets needed for commissioning were installed in both recirculation arcs along with
the trim quads.
5) The large recirculation magnet needed for the second arc (GY) arrived this week along with two more reverse
     bend dipoles.  Only 3 more magnets are outstanding from the vendor in order to complete the second arc and
the upgrade beam transport system.  We expect these in approximately two weeks.
6) The wiring was completed for the high reflector optical cavity.
7)  The RF supply which misbehaved during SRF commissioning of the injector last week was fixed.

On Thursday, Jan. 16th, Fred Dylla and Henry Helvajian (from Aerospace Corporation) visited the Air Force
Research Laboratory (AFRL) for a seminar and discussions on the UV FEL and UV Microfabrication Station.
Our AF Program Manager, John Eric hosted the visit and arranged for discussions with Barry Farmer, the Chief
Scientist of the AFRL Materials and Manufacturing Directorate, and with Barbara Wilson, the Chief Technologist
of AFRL.

WBS 4 (Injector):
We degassed the cesiators, heat cleaned and prepared a cathode and mapped the QE.  Measured QE's were
more than sufficient for initial operation (QE~0.5%).  The combined running and conditioning resistors were
installed and the HVPS system made operational.  The Gun was HV processed to 50 kV as of Thursday night.
We noted that we were getting occasional field emission spikes at these low voltages.  Upon inspection of the
GaAs cathode on Friday morning we noted that it suffered a small crack near the perimeter, possibly as a result of
interaction with the door mechanism.  The door mechanism is safely in the open position now with non interference
with the cathode.  Given the progress we have made so far with the remainder of the injector and linac, we decided
to replace only the GaAs cathode and do a bakeout of only the gun assembly in order to proceed with beam
commissioning.  We will defer a full bakeout of the whole system only when we are limited by system performance.

WBS 6 (RF):
Quarter HVPS - The HVPS for Quarter Unit cavity #4 was repaired and operated this week.  The firing circuit
boards for the SCR controller were returned to the vendor for testing and repair as required.  New spares were
ordered as was several transformers to facilitate any future repairs in house.  Both Quarter units were operated into
the their cavities, but cryo instabilities frustrated any significant RF testing.

Buncher - The Buncher has been reassembled and tested.  It is operational.

Zone 2 - A faulty RF Crate was replaced.  The zone is fully operational.

Zone 3 - The firing circuit boards from this zone's SCR controller were removed to repair the controller in Quarter
#4.  Replacement boards are on order.

Zone 4 - In order to provide more drive to this zone's klystrons, the input attenuators to the new 5 W preamplifiers
were reduced from 6 dB to 3 dB.  We were not yet able to test this change due to cryo instabilities last night.

WBS 8 (Instrumentation):
The week was spent preparing for gun commissioning and beam ops.  Most of the systems are now fully
operational and ready for beam.

The Ops screens for the trim magnets and box supplies from 0F to 5F have been completed and are located under
the magnet commander page.  Rack termination's for the 2nd Arc, Wiggler, dispersion and injection shunts
continue.  Drive laser LVDT's and Pico motor termination's in the Injector have been completed.  Injector video
wiring was pulled back into the control room from the vault and re-installed.  Checkout of the High Reflector
LVDT and Stepper Motor systems are in progress.  The second stepper motor chassis was completed.

The 4-channel BPMs in the 1F and 2F region have been connected and signal checked.  Drawings for the HVPS
Interlock Box are complete as well as several miscellaneous test boxes. The new analog control card for the gun
power supply is here and being tested.  The synchrotron light monitor attenuator part also came in and will begun to
be assembled next week.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
• No progress
Diversion of the beam after the Wiggler.
• The vacuum chamber was installed in the GW magnet.
Optical Chicane Dipoles (GW)
• Magnet Test will do testing on these UV Line magnets after the IR Magnets are fully tested.
Injector Dipoles (DU/DV)
• No action
Arc 180 Degree Dipoles (GY)
•  The first GY Magnet was received on Monday.  Don Bullard started putting together the semicircular-tracked
    magnet measurement apparatus developed on the IR Demo to measure the magnet’s performance.  Don became

    available because his gun duties have stopped now that high voltage processing has started.  He may be pulled at

    any moment if more work is necessary on the Gun.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
•  Data was taken on the GX over an area with double the width of previous measurements.  We were able to take

    the data in the standard IR configuration as well as the skewed UV configuration. David Douglas is continuing to
    have Chris Tennant, work toward analyzing the field integrals over beam paths using the graphics program
•  Measurements of the GX had to be restarted several times because they were interrupted as the power supply
    turned off, due to phase mismatch, three hours into the 4-hour measurement.  Simon Wood got the power
    going again.  Hopefully the fix is permanent.
•  The remaining three GX magnets were modified in the shop to remove the post in the return leg and allow the
    wedge of the magnet to shorten its effect.
Trim Quad (QT)
•  The trim coils to create the GC vertical corrector function within the four trim quads are half complete at Milhous
    Control.  They project completion on January 22.
Sextupole (SF)
•  The Sextupole Light magnets were mounted in their positions within the enclosure.
Octupole (OT)
•  The procurement process started up again with a "No Funds" requisition to get bids so that as soon as funds are
    available, the fabrication can begin.
Beam Line and Vacuum
•  Our shop continued working on the last chamber in Optical Chicane.  They forecast completion on the 24th of
•  Two (of four) qualified chambers of the Optical Chicane were installed in the respective GW magnets in the
•  The two 180* Bend chambers await the receipt of the diagnostic windows from Ceramaseal.  The latest
    delivery date is early February.
•  The Vacuum Group connected girders in the regions before and after the second arc in their drive to button up
•  With the help of visiting Scientist, Harvey Rutt we were able to install a number of corrector dipoles in the
    and the remaining portions of the machine.  This activity was stopped earlier because the installers (Magnet
    Measurement folks) couldn’t be distracted from their measurement duties.

WBS 11 (Optics):
Wiring for the HR optical cavity assembly is essentially complete.  Connection to the controls system begins next
week.  We will receive components to bypass cooling water for deformable mirror assemblies we aren't ready to
install.  After we complete that work we should be able to put that assembly under vacuum and continue
commissioning.  As for the OC assembly, we received the molybdenum required to fabricate the edge-cooled
mirror holder.  Since fab and subsequent plating will take some time, we have fabricated a simple aluminum adapter
ring that we will us for the first light activities.  Fabricated parts for the diagnostic viewer/laser safety shutter and the
insertable mirrors are arriving and we are on schedule to receive all parts the third week of January.  A long lead
item, the custom air cylinders that actuate the insertable mirrors, arrived this week.

Other commissioning activities:
We finished realigning the laser through the optics in the clean room this week.  We reinstalled the picomotor
controller, and discovered the "c" axis channel blew out again.  Apparently the vendor did not find the root cause of
the problem.  We have a temporary solution that will get us through the next few weeks of commissioning.  All
picomotors in the drive laser OTS check out, as do the LVDTs in the last mirror can.  Thanks to I&C for getting
the cabling installed and terminated.  We plan to align the beam through it early next week.